OpenStudy (anonymous):

given f(x)=-x^2+2x, find f(-3) will medal and fan :3

Nnesha (nnesha):

f(-3) that means x =-3 so substitute x for -3 into f(x)=-x^2+2x

OpenStudy (anonymous):

f(-3)=-3x^2+2(3)

OpenStudy (anonymous):

i didn't mean that x after 3...

Nnesha (nnesha):

alright it should be -(-3)^2+2(-3) now solve this

OpenStudy (anonymous):

-(-9)+-6

Nnesha (nnesha):

when you take even power of negative number you will always get positive answer

OpenStudy (anonymous):

so 9+6?

Nnesha (nnesha):

\[\huge\rm -\color{ReD}{x}^2+2\color{ReD}{x}\]this is the function replace x with -3 \[\huge\rm -(\color{ReD}{-3})^2+2(\color{ReD}{-3})\] 2 times -3 = ?

OpenStudy (anonymous):

-6

Nnesha (nnesha):

yes right \[-(-3)^2-6\] solve

OpenStudy (anonymous):

9-6=3

Nnesha (nnesha):

(-3)^2 = 9 there is a negative sign front of the parentheses

OpenStudy (anonymous):

but you said two negatives make a positive????

Nnesha (nnesha):

yes right so (-3)^2 = -3 times -3= 9

OpenStudy (anonymous):

i got 9 and i plugged it into the equation and that's how i got 3

Nnesha (nnesha):

\[\huge\rm\color{blue}{ -}\color{ReD}{x}^2+2\color{ReD}{x}\]this is the function replace x with -3 \[\huge\rm \color{blue}{-}(\color{ReD}{-3})^2+2(\color{ReD}{-3})\] \[\huge\rm \color{blue}{-}(\color{ReD}{9})-6\] (-3)^2 =9 but i'm saying blue sign would stay there

Nnesha (nnesha):